# Mapping Trustworthiness in Large Language Models: A Bibliometric Analysis Bridging Theory to Practice

### v1

@@ 4.6. Thematic Evolution
Regarding AI ethics and responsible AI, terms used in this period, they are discussed through theoretical principles (Jobin et al., 2019), (Floridi and Cowls, 2022), (Mittelstadt, 2019), (Vakkuri et al., 2021) and the urge to translate these principles into practice (Morley et al., 2019b), (Vakkuri et al., 2020), (de Cerqueira et al., 2022), (Siqueira De Cerqueira et al., 2021).
As AI systems – especially LLMs – became more advanced, concerns over trust, transparency, and accountability became even more pronounced. In 2025-2025, Figure 4 reveals a diversification of research topics, branching into AI in education, blockchain, trustworthiness, creativity, chatbots, and disinformation.
The rise of trustworthiness comes mainly from the term responsible AI, this demonstrates how the discourse has shifted from traditional responsible AI principles in AI ethics to what is trustworthiness in generative AI. This progression highlights the tension between innovation and ethical oversight, reinforcing the need for concrete trustworthiness frameworks rather than abstract trustworthiness taxonomies.
Similarly, the focus on AI ethics has shifted to concerns about disinformation, reflecting the fact that LLMs are rapidly replacing traditional search engines and are primarily used as chatbots, particularly in question-answering (QA) settings. As users increasingly rely on LLMs rather than traditional search engines to retrieve information, the risk of misinformation and manipulated content has become a central issue. This typical type of use of LLMs suggests why discussions of AI ethics have evolved to disinformation.
Figure 4 aligns with the bibliometric mapping approach of this study, revealing how AI ethics has been rebranded from abstract guiding principles into abstract trustworthiness taxonomies for LLMs.
Researchers that have never delved on Artificial Intelligence suddenly started researching about their ethical issues but in the generative AI ChatGPT era, forcing them to come up with new old terminologies that have been around for almost a decade. This sudden interest is, in the opinion of the first author of this paper, hindering the actual advancements towards more ethical AI, as it is masking the debate towards a ”new” set of theoretical guidelines as an enforced contribution, but not providing practical guidance for the developers, the ones actual responsible for the development of ethical AI.
In other words, this perceived rebranding of responsible AI and AI ethics to trustworthiness and disinformation is for the most part failing in proposing useful unified terminologies, frameworks, or practical guidance for developers to operationalize principles.
There is a growing need for enforceable guidelines beyond non binding ethical commitments. Even more evident after big tech companies such as Google withdrawing its AI ethical principles in mass surveillance and AI weapons (Hooker and Vallance, 2025). Consequently, there is an urge for actual hard laws that can govern the development and use of these systems, that can move beyond the EU AI Act (Sillberg et al., 2024), due to the fact that it only propose risk level assessments, and links to the AI HLEG (High-Level Expert Group on Artificial Intelligence, 2019). The later has a set of trustworthy AI principles, however, it is non binding. Despite the fact that a binding document links to a non binding one, does not automatically transform the AI HLEG into hard law. On top of that, an unified set of practical guidance are still lacking, after several years of debate – at least since 2019 (Morley et al., 2019b).
Despite this progress, fragmentation remains, with multiple authors competing with definitions of LLM trustworthiness rather than a single, unified framework.
The continued proliferation of non binding framework proposals and terminologies by singular researchers highlights the urgency of bridging the gap between theoretical AI ethics and its real-world implementation through unified frameworks and legislations.
Figure 4. Thematic Evolution: 2019-2025.
The word cloud in Figure 5 provides a visual representation of the most common keywords in the literature related to LLM, genAI and its trust and ethical implications. The size of each keyword indicates its frequency, with larger words appearing more frequently in the reviewed studies.
Among the most prominent terms, ”ChatGPT”, ”trust”, ”AI”, ”technology”, and ”information” are the most frequently mentioned, indicating that discussions around trust and AI adoption play a crucial role in ongoing research. It also shows that researchers are mainly making observations based on the ChatGPT platform, which is certainly the main LLM used in academia.
In addition, ”health” and ”education” emerge as recurring themes, reflecting concerns about the impact of AI in different domains, and broadening the debates about the responsible use of AI and its impact on different domains, particularly education and healthcare.
The word cloud also highlights other frequently cited themes, such as ”performance,” ”user acceptance,” ”privacy,” ”ethics,” ”impact,” ”challenges,” ”bias,” ”intention,” ”risk,” and ”opportunities.” These recurring themes highlight the growing emphasis on trustworthy AI, with concerns such as bias, privacy and ethics closely linked to user acceptance. At the same time, the literature recognises both challenges and opportunities, emphasising the need for responsible AI adoption while exploring its potential benefits.
Overall, Figure 5 illustrates the multifaceted nature of ChatGPT research, encompassing technical advances, ethical considerations, societal impacts, and real-world applications, particularly in areas such as trust in AI, automation, education, and user interaction. The findings reinforce the importance of ensuring ethical implementation of AI, transparent models and responsible innovation in this rapidly evolving field.
Figure 5. Word cloud.
4.7. Co-Occurrence Network by Keyword Plus
The co-occurrence network visualization in Figure 6 highlights the central role of ”artificial intelligence”, with strong associations to key themes like ”ChatGPT”, ”chatbot”, ”large language models”, related to the technologies themselves, and to ”medicine”, ”science”, ”education”, ”health”, relating to domains of interest; and ”ethics” and ”bias” emphisizing the ethical concerns. This suggests that AI research is deeply interconnected with societal and ethical challenges, particularly in LLMs and their impact on various domains.
On the other hand, the blue cluster focuses on the concept of ”trust”, linking it to ”acceptance”, ”user acceptance”, ”perceptions”, ”adoption”, ”quality” and ”privacy”. This indicates a growing interest in AI trustworthiness, highlighting it as in terms of user acceptance. That is, instead of having previous set of AI ethical guidelines governing AI use, development and deployment in a more holistic way, this suggests a shift in a more simplistic view of trust in AI is simply gauging the acceptance of a LLM by the user.
The co-occurrence network in Figure 6 highlights the the divide between ethical AI development and its trust/trustworthiness implications.
Figure 6. Co-Occurrence Network.
The clustering by coupling analysis presented in Figure 7 categorizes research themes based on author coupling, using references as a measure. The x-axis (Centrality) represents the extent to which a topic is interconnected with others, while the y-axis (Impact) reflects the influence of each cluster based on global citation scores. The clusters highlight key themes in AI research, with a strong focus on artificial intelligence, ChatGPT, trust, and performance. Notably, the red cluster (AI, trust) has high confidence values, suggesting a strong association between AI discussions and trustworthiness concerns.
The green cluster, positioned near the center, contains artificial intelligence, ChatGPT, and performance, indicating a balance between centrality and impact. This suggests that discussions about ChatGPT’s capabilities and AI performance are well-integrated within broader AI research. The top-right cluster, though smaller, highlights decision-making and AI systems, implying its emerging importance in AI governance. The analysis suggests a growing intersection between AI ethics, trust, and performance, reinforcing the need for responsible AI development and evaluation frameworks.
Our main interest resides in the red cluster, as it discusses trust. Table 4 shows the most important authors ranked by its relevance within the red cluster.
Overall, the five studies highlight the rapid adoption and diverse applications of large language models (LLMs) across multiple domains, while also highlighting their ethical, practical, and evaluative challenges. Martinez-Maldonado et al. systematically review the use of LLMs in education and identify nine key applications, such as grading, content generation and feedback, while cautioning about replicability, transparency and ethical concerns. Park et al. extend this discussion by investigating the impact of misinformation generated by ChatGPT on tourism decision-making, demonstrating how trust in AI can be easily manipulated when incorrect information is prominently presented. Similarly, Sallam examines AI in healthcare, highlighting its potential to improve medical research, education and clinical workflows, but also highlighting risks such as bias, hallucination, plagiarism and cybersecurity threats.
Wang et al. provide a meta-analysis of AI evaluation methods and argue for a more structured approach to assessing their performance, ethical considerations and real-world impact in different domains. This ties in with Herrera et al.’s work on Explainable AI (XAI), which highlights the urgent need for transparency in AI decision-making and proposes 28 open problems for achieving better interpretability and accountability in AI systems. Taken together, these studies show that while LLMs have transformative potential, their responsible use requires robust evaluation methods, ethical guidelines and accountability.
@@ 4.7. Co-Occurrence Network by Keyword Plus
1 Martinez-Maldonado R Practical and ethical challenges of large language models in education: A systematic scoping review (Yan et al., 2023)
@@
2 Park J When ChatGPT Gives Incorrect Answers: The Impact of Inaccurate Information by Generative AI on Tourism Decision-Making (Kim et al., 2023)
@@
3 Sallam M ChatGPT Utility in Healthcare Education, Research, and Practice: Systematic Review on the Promising Perspectives and Valid Concerns (Sallam, 2023)
4 Wang C A Survey on Evaluation of Large Language Models (Chang et al., 2023)
@@
5 Herrera F Explainable Artificial Intelligence (XAI) 2.0: A manifesto of open challenges and interdisciplinary research directions (Longo et al., 2024)
Table 4. Top Cited Authors within ’trust’ Cluster and Their Works
Figure 7. Clustering by Coupling.
4.8. The LLM trust/trustworthiness definitions proposed by different authors in the literature
@@
In Table LABEL:tab:trust_definitions, we provide a structured overview of various definitions and conceptualizations of trust and trustworthiness in AI, as discussed in the literature. This compilation allows for a comparative analysis of how trust in AI is framed across different domains, emphasizing dimensions such as ability, reliability, integrity, explainability and ethical alignment. From the initial 68 analysed documents, 18 documents are explicitly or implicitly providing useful information to the task of framing what is trust/trustworthiness in LLM or genAI. That is, 26.47% against 73.53% of documents that are not providing useful information for this task. Other documents were not taken into consideration because they are either closed access, editorial, on the use of LLM to improve trust, in a different domain e.g., radiology, measuring trust level of a certain group when using certain LLM system, or simply with a different focus. Interesting to note that 5 of the discarded papers were related to fake news, misinformation and deepfake. The complete analysis is available on Zenodo.
@@
Trust is a multidimensional construct that encompasses several key components, including benevolence, competence, and integrity (Mayer et al., 1995). Mayer et al.’s work entitled ”An Integrative Model of Organizational Trust” was published in 1995 in a Management venue, and has been adapted in several contexts, one of them being in trust in AI systems. We found a tendency in reviewed studies in utilizing Mayer et al.’s work to build on definitions for what is trustworthiness in LLM – 33,33%, one third, of definitions (6/18 entries: 5 (Wang et al., 2024a), 6 (Ganapati and Desouza, 2024), 8 (Lacity et al., 2024), 12 (Ressel et al., 2024), 15 (Jia et al.(2025), Jasper), 16 (Martell et al., 2024)) explicitly draw from Mayer et al.’s (Mayer et al., 1995) organizational trust framework. Its core elements – ABI+ factors of perceived trustworthiness – towards AI, are (Martell et al., 2024):
  1. (1)
@@
Trust in AI, particularly in situations of uncertainty and vulnerability, is described as the belief that an AI agent will support an individual in achieving his or her goals (Wang et al., 2024a). Trust is fundamentally linked to vulnerability and uncertainty, as highlighted by multiple perspectives. Hoy and Tschannen-Moran (Hoy and Tschannen-Moran, 1999) define trust as a willingness to be vulnerable based on confidence in another party’s benevolence, reliability, competence, honesty, and openness, a concept applied to AI-mediated teacher-student relationships. Mayer et al. (Mayer et al., 1995) reinforce this by framing trust as a positive expectation of another’s intentions and behavior, emphasizing the risk involved in relying on an entity. Similarly, Lee and See (Lee and See, 2004) describe trust as the belief that an agent will assist in achieving one’s goals despite uncertainty and vulnerability, underscoring that trust in AI systems depends not just on their performance but also on users’ confidence in their intent and reliability.
In this context, users assume the vulnerable position of relying on genAI models or proprietary platforms with the expectation that this risk will be beneficial to them, i.e. that it will help them achieve their goals despite possible misinformation or bias (Ressel et al., 2024). In the paper ”Mitigative Strategies for Recovering From Large Language Model Trust Violations” (Martell et al., 2024), the authors present three different levels of trust calibration, as shown in Table 5.
Table 5. Types of Trust Calibration in Human-AI Interaction (Martell et al., 2024)
@@
Appropriate (Calibrated) Trust: Alignment between an individual’s perceived trust level and the AI system’s actual capabilities and performance. (Yang et al., 2020)
Over-Trust: Occurs when users: Perceive the AI as more capable than it truly is Treat the technology as a fully competent teammate Over-rely on AI outputs despite known limitations (Aroyo et al., 2021)
Under-Trust: Occurs when users: Underestimate the AI system’s capabilities Over-rely on their own judgment Dismiss potentially valuable AI contributions (Hoff and Bashir, 2014)
@@
The more familiar we are with a tool, the more we tend to trust them (Califano et al., 2024). Other authors state that trust in AI-LLM can be translated as the degree to which users and institutions are confident that the data used by, and the information generated by, AI-LLM is impartial, i.e. unbiased (Jenks, 2024).
Two more trust related dimensions are presented in Table LABEL:tab:trust_definitions: cognitive trust (rational evaluation of a system’s reliability and accuracy) and affective trust (emotional confidence in its benevolence and alignment with user interests) (Zhang et al., 2014). Cognitive trust, anchored in perceived credibility, facilitates affective trust through sustained reliance, as users transfer trust from the platform’s technical competence to its perceived ethical intent (Zhang et al., 2014; Wang et al., 2016). This interdependence highlights how platform reliability shapes users’ perceptions of AI-generated content, with technical accuracy (’reliable’) and empathetic design (’credible’) jointly calibrating holistic trust.
What are the trade-offs of using or adapting a 30 year old organizational trust framework to the AI era? First, as a benefit, it is a well established set of trust taxonomy, that has been widely accepted in the academia. However, it was devised way before wider adoption of different AI technologies, in special LLMs. Similarly to AI ethics, where soft law recommendations did not pose serious changes, should trustworthiness in LLM discussion be on hard laws that can unify the terms and recommendations?
In (Jenks, 2024), the authors state three principles of AI-LLM: fair, unbiased, and ethical. On top of that, they relate these principles with the work done within the European Commission on AI ”Building Trust in Human-Centric Artificial Intelligence” (Commission, 2019), which states that trustworthy AI should not only be lawful, ethical, and robust, but also give humans the ability to have control over its use, draw from and provide reliable data, offer security and privacy to its users, and be transparent, accountable, and inclusive. The efforts of the European Commision on AI is better translated in the EU AI Act (Sillberg et al., 2024; Commission, 2023), where a risk level assessment for AI systems is provided. We briefly describe the EU AI Act’s risk level in Table 6. From this Table, we pinpoint that most of genAI and LLM generated content falls under the ”Limited Risk” level, where there is a need for Transparency requirements, highlighting once again the importance of this trustworthy principle.
Risk Level Description
Unacceptable Risk AI systems that are prohibited due to threats to fundamental rights (e.g., social scoring, manipulation).
High Risk AI systems that impact safety or fundamental rights (e.g., biometric ID, credit scoring, law enforcement).
Limited Risk AI systems with transparency requirements (e.g., chatbots, AI-generated content).
Minimal Risk AI systems with no specific regulations (e.g., spam filters, AI-driven recommendations).
Table 6. Risk Levels in the EU AI Act
In addition, within the EU AI Act there is reference to a non-binding document, also provided by the European Comission, that aims to provide a set of trustworthy principles in AI. This non-binding document is the 2019 Ethics Guidelines for Trustworthy AI developed by the AI High-Level Expert Group (HLEG) (High-Level Expert Group on Artificial Intelligence, 2019). We shortly describe its principles in Table 7. This Table should complement our Table 9, as there are many overlapping terms, such as Transparency (Transparency) and Accountability (Responsible).
Principle Description
Table 7. AI HLEG Principles for Trustworthy AI
Finally, in (Baeza-Yates, 2024), there is a critic towards the use of the terms trustworthy and ethical AI, as they can be used to humanize, i.e., anthropomorphise AI. This critic is valid insofar every Information System artefact, such as LLM and AI systems, are devised by humans, and they are programmatically designed to fulfil a set of subjective or objective desires implied by the developers or organizations they are inserted.
Table 8. Trustworthiness Definitions in AI
1 Foundation Models (FMs) and Large Language Models (LLMs) demonstrate strong generalization capabilities and adaptability to specific tasks through fine-tuning or in-context learning. However, they currently lack the predictability and robustness required for deployment ’as is’ in adversarial environments. In this context, trustworthiness is characterized by the ability to be directly deployed in real-world scenarios while maintaining resilience against adversarial conditions. Trustworthy, resilient, and interpretable artificial intelligence (AI) (Jha et al., 2023) -
2 To improve trust (in these models) is to understand the behavior of deep neural networks for Information Retrieval (IR). Explainable AI (Lucchese et al., 2023) -
3 Not explicitly, but states that: LLM’s ability in providing reliable (S&P) advice. Trusted, Reliable (Chen et al., 2023) -
4 The authors argue that the perceived trustworthiness of GAN-generated content depends on several factors, including the realism of the content, the context in which it is presented, and the pre-existing trust individuals have in the sources of information. More specifically in the case of fake news and misinformation. The authors suggest that addressing the ethical and political dimensions of GANs is crucial for maintaining trust in democratic societies. Perceived trustworthiness, disinformation concerns (Carnevale et al., 2023) -
5 Trust in AI is described as the belief that an AI agent will support an individual in achieving their goals, particularly in situations marked by uncertainty and vulnerability. Trust plays a crucial role in user interaction with AI, as its absence can lead to reluctance in adopting AI tools, even when they demonstrate superior performance. Conversely, excessive trust in AI, especially in high-stakes domains like software engineering, can cause users to overlook potential errors or risks. The study explores strategies for fostering appropriate trust in generative AI tools, such as GitHub Copilot, to enhance their usability and reliability. Ability, Benevolence, Integrity (Wang et al., 2024a) (Lee and See, 2004) (Liao and Sundar, 2022) (Vereschak et al., 2021) (Commission, 2019) (Das and Rad, 2020) (Boubin et al., 2017) (O’Connor et al., 2019) (Pearce et al., 2022) (Perry et al., 2023) (Mayer et al., 1995)
6 Trust in AI comprises three key components: ability (ensuring reliable performance), integrity (aligning with ethical and social norms), and benevolence (acting in society’s best interest). Trustworthiness is reinforced by verifying AI’s reliability, ensuring ethical alignment, and strengthening oversight through public scrutiny and regulation, emphasizing governance over technical solutions. Ability, integrity, benevolence, verifiability (Ganapati and Desouza, 2024) (Desouza and Dawson, 2023) (Mayer et al., 1995)
7 While not formally defining trustworthiness, it conceptualizes it as the ability of AI systems to avoid generating false or misleading responses by accurately determining when they lack sufficient evidence to answer a given question. Reliable, factual (Łajewska and Balog, 2024) -
8 Trust in IT artifacts is defined by their quality, reliability, accuracy, and security. While organizational trust frameworks, such as Mayer et al. (Mayer et al., 1995), are widely used in IS research, applying concepts like benevolence to IT artifacts risks anthropomorphizing technology. Ability/competence, integrity, and benevolence (Lacity et al., 2024) (Mayer et al., 1995) (Porra et al., 2019) (Belanger and Crossler, 2019) (Kim et al., 2009) (Lowry et al., 2013)
9 Not explicitly, but it describes trustworthy AI in terms of consistency, reliability, explainability, and safety. Consistency, reliability, explainability, and safety (Gaur and Sheth, 2024) -
10 It follows a definition of trust in education, which states that: ”Trust is an individual’s or group’s willingness to be vulnerable to another party based on the confidence that the latter party is benevolent, reliable, competent, honest, and open.” This definition is applied to teacher-student relationships to explore how trust is built and eroded in an AI-mediated assessment environment. Transparency, reliability, honesty, competence, integrity, fairness, benevolence, openness, trust-building (Luo, Jess) (Hoy and Tschannen-Moran, 1999)
11 Believing in the credibility or accuracy of a statement. Credibility, AI alignment (Buchanan and Hickman, 2024) -
12 Trust is characterized by a willingness to be vulnerable, accompanied by a positive expectation of the other party’s intentions and behavior. In the context of AI and LLMs, real-world incidents have urged the need for translating high-level trustworthiness rhetoric into domain-specific practices. This raises the question of how trust should be defined in fields such as the insurance industry and whether regulatory bodies should actively promote trust in AI systems. AI ethics, AI trustworthy (Ressel et al., 2024) (Mayer et al., 1995)
13 While trust in AI-generated content is not explicitly defined, the authors suggest that trust can be enhanced by providing indicators of familiarity, validation, or endorsements from experts. However, they also note that people tend to be more cautious when AI-generated content is explicitly labeled. Transparency, disclosure, validation, familiarity, experience, task suitability, credibility (Califano et al., 2024) -
14 Trust in AI-LLM is defined as confidence in the impartiality of its data and generated information. Personalization plays a crucial role in fostering trust in AI-LLM, particularly in intercultural communication. According to the European Commission (2024), trustworthy AI should not only be lawful, ethical, and robust but also empower users with control over its use, ensure data reliability, security, and privacy, and uphold transparency, accountability, and inclusivity. Ultimately, trust in AI-LLM depends on the perception that its outputs are reliable and unbiased, with trustworthiness requiring ethical, inclusive, and accountable AI-generated knowledge. Confidence, fair, unbiased, ethical, lawful, robust, human oversight, transparency, accountability, security, privacy (Jenks, 2024) (Cook, 2001) (Araujo et al., 2020) (Commission, 2025)
15 Trust is a multidimensional construct that encompasses several key components, including benevolence, competence, and integrity. Research suggests that users tend to trust AI-generated content less when they are aware of its artificial origin, a trend observed in multiple studies. Credibility, Transparency (Jia et al.(2025), Jasper) (Mayer et al., 1995)
16 Trust is defined as the belief that an agent will assist in achieving an individual’s goals in situations marked by uncertainty and vulnerability. Appropriate trust, or calibrated trust, occurs when a person’s trust in a system aligns with its actual performance. Over-trust arises when individuals overestimate AI capabilities, while under-trust occurs when users undervalue AI and rely too heavily on themselves. The ABI+ framework outlines four key dimensions of trust: (1) Ability – AI’s capacity to perform tasks effectively, (2) Benevolence – AI’s intent to act in users’ best interests, (3) Integrity – adherence to ethical standards, and (4) Predictability – consistency in AI responses. This highlights the importance of ensuring AI systems maintain accuracy from early deployment stages, as initial errors can have long-term consequences on user trust calibration and later adoption. Transparency, calibrated trust, reliability (Martell et al., 2024) (Mayer et al., 1995) (Lee and See, 2004) (Yang et al., 2020) (Aroyo et al., 2021) (Hoff and Bashir, 2014) (Dietz and Den Hartog, 2006) (Toreini et al., 2020)
17 Trust in generative AI can be divided into cognitive trust and affective trust. Cognitive trust is based on a user’s evaluation of the AI platform’s ability to deliver accurate and reliable information, while affective trust stems from the user’s belief that the AI system considers their interests and well-being. Reliable, credible (Zhou and Lu, 2024) (Zhang et al., 2014) (Wang et al., 2016) (Stewart, 2003)
18 The author actually criticizes the term trustworthy and ethical AI because they can humanize AI. Responsible AI (Baeza-Yates, 2024) -
In Table 9, we present the frequency of key terms associated with trustworthiness in AI, reflecting the most commonly emphasized aspects in the literature. Transparency and Explainability appear as the first and second most frequently mentioned characteristics, respectively, highlighting how they are enabling principles, i.e., without transparency and explainability it is impossible for other trustworthiness or ethical principles to exist (Siqueira De Cerqueira et al., 2021). Reliability is related to Ability and Predicability in the ABI+ framework (Mayer et al., 1995). Furthermore, Ethics, Privacy, and Fairness highlight concerns related to Integrity, i.e., adherence to ethical standards. This Table 9 provides a quantitative perspective on the many trustworthiness aspects identified from the literature.

### v2

@@ IV Results
By 2025, the thematic scope diversified further, branching into AI in education, blockchain, trustworthiness, creativity, chatbots, and disinformation. This evolution demonstrates the increasing complexity and multidisciplinary nature of the LLM trustworthiness debate.
The word cloud in Figure 4 represents the most frequently occurring keywords in the LLM trustworthiness and ethics literature. The size of each word corresponds to its relative frequency in the data set.
Notably, the terms ‘ChatGPT’, ‘trust’, ‘AI’, ‘technology’ and ‘information’ appear as the most dominant. Other common terms include ‘health’, ‘education’, ‘performance’, ‘impact’ and ‘user acceptance’. The presence of domain-specific terms such as ‘health’ and ‘education’ indicates relevant domains where LLMs are applied. In addition, several ethical related keywords – such as ‘privacy’, ‘ethics’, ‘bias’ and ‘challenges’ – indicates the researchers interest in this growing area.
Figure 4: Most frequently occurring keywords in LLM trust and ethics literature (2019–2025)
We analyse studies that reference the same arrangement of cited articles in order to understand the bibliographic coupling. Through this analysis it is possible to cluster authors and identify possible themes. In Figure 5 four clusters are visible, highlighting the importance in academia of linking AI to 1) trust, 2) performance, 3) recommendations and 4) decision-making. Also noticeable is the appearance of ChatGPT in two clusters, in relation to performance and recommendations. Our main interest is focused on the red cluster, as it deals with trust. In this sense, we analyse studies from the five most prominent authors within this theme. Table I shows the authors ranked by their relevance within the red cluster.
Figure 5: Clustering by Coupling.
TABLE I: Top Cited Authors within ’trust’ Cluster and Their Works
@@
1 Martinez-Maldonado R Practical and ethical challenges of large language models in education: A systematic scoping review [22]
@@
2 Park J When ChatGPT Gives Incorrect Answers: The Impact of Inaccurate Information by Generative AI on Tourism Decision-Making [23]
@@
3 Sallam M ChatGPT Utility in Healthcare Education, Research, and Practice: Systematic Review on the Promising Perspectives and Valid Concerns [21]
4 Wang C A Survey on Evaluation of Large Language Models [24]
@@
5 Herrera F Explainable Artificial Intelligence (XAI) 2.0: A manifesto of open challenges and interdisciplinary research directions [25]
To address Research Questions 2 and 3, we manually analyzed 68 articles in which the term “trust*” appeared in the abstract.
TABLE II: Trustworthiness Definitions in AI
1 LLMs demonstrate strong generalization capabilities and adaptability to specific tasks through fine-tuning or in-context learning. However, they currently lack the predictability and robustness required for deployment ’as is’ in adversarial environments. In this context, trustworthiness is characterized by the ability to be directly deployed in real-world scenarios while maintaining resilience against adversarial conditions. Trustworthy, resilient, and interpretable artificial intelligence (AI) [26] -
2 To improve trust is to understand the behaviour of deep neural networks for Information Retrieval. Explainable AI [27] -
3 Not explicitly, but states that: LLM’s ability in providing reliable security and privacy advice. Trusted, Reliable [28] -
4 The perceived trustworthiness depends on several factors: the realism of the content, the context in which it is presented, and the pre-existing trust individuals have in the sources of information. Notably relevant to misinformation contexts. Addressing the ethical and political dimensions is crucial for maintaining trust in democratic societies. Perceived trustworthiness, disinformation concerns [29] -
5 Trust in AI is described as the belief that an AI agent will support an individual in achieving their goals, particularly in situations marked by uncertainty and vulnerability. Trust plays a crucial role in user interaction with AI, as its absence can lead to reluctance in adopting AI tools, even when they demonstrate superior performance. Conversely, excessive trust in AI, especially in high-stakes domains like software engineering, can cause users to overlook potential errors or risks. The study explores strategies for fostering appropriate trust in generative AI tools, such as GitHub Copilot, to enhance their usability and reliability. Ability, Benevolence, Integrity [30] [31, 32, 33, 34, 35, 36, 37, 38, 39, 10]
6 Trust involves ability (providing reliable performance), integrity (aligning with ethical and social norms), and benevolence (acting in society’s best interest). Trustworthiness is reinforced by verifying AI’s reliability, yielding ethical alignment, and strengthening oversight through public supervision and regulation, emphasizing governance over technical solutions. Ability, integrity, benevolence, verifiability [40] [41, 10]
7 Not formally, but conceptualizes it as the ability of AI systems to avoid generating false or misleading responses by accurately determining when they lack sufficient evidence to answer a given question. Reliable, factual [42] -
8 Trust in IT artifacts is defined by their quality, reliability, accuracy, and security. While organizational trust frameworks, such as Mayer et al. [10], are widely used in IS research, applying concepts like benevolence to IT artifacts risks anthropomorphizing technology. Ability/competence, integrity, and benevolence [43] [10, 44, 45, 46, 47]
9 Not explicitly, but it describes trustworthy AI in terms of consistency, reliability, explainability, and safety. Consistency, reliability, explainability, and safety [48] -
10 “Trust is an individual’s or group’s willingness to be vulnerable to another party based on the confidence that the latter party is benevolent, reliable, competent, honest, and open.” This definition is applied to teacher-student relationships to explore how trust is built and eroded in an AI-mediated assessment environment. Transparency, reliability, honesty, competence, integrity, fairness, benevolence, openness, trust-building [49] [50]
11 Believing in the credibility or accuracy of a statement. Credibility, AI alignment [51] -
12 A willingness to be vulnerable, accompanied by a positive expectation of the other party’s intentions and behavior. Real-world AI incidents have urged the need for translating high-level trustworthiness rhetoric into domain-specific practices. This raises the question of how trust should be defined in fields such as the insurance industry and whether regulatory bodies should actively promote trust in AI systems. AI ethics, AI trustworthy [52] [10]
13 Not explicitly, but authors suggest that trust can be enhanced by providing indicators of familiarity, validation, or endorsements from experts. However, people tend to be more cautious when AI-generated content is explicitly labeled. Transparency, disclosure, validation, familiarity, experience, task suitability, credibility [53] -
14 Trust is defined as possessing a confidence in something or someone. While personalization is key to fostering trust, the black-box problem – lack of transparency and explainability – hinders trust. Trustworthy AI should not only be lawful, ethical, and robust, but also give humans the ability to have control over its use, draw from and provide reliable data, offer security and privacy to its users, and be transparent, accountable, and inclusive. Trust in AI-LLM means that individuals and institutions are confident that the data used by, and information generated in, AI-LLM are impartial. Confidence, fair, unbiased, ethical, lawful, robust, human oversight, transparency, accountability, security, privacy [54] [55, 56, 57]
15 Trust is a multidimensional construct that encompasses several key components, including benevolence, competence, and integrity. Users tend to trust AI-generated content less when they are aware of its origin. Credibility, Transparency [58] [10]
16 Trust is defined as the belief that an agent will assist in achieving an individual’s goals in situations marked by uncertainty and vulnerability. Addressing trust is fundamental from early deployment stages, as initial errors can have long-term consequences on user trust calibration and later adoption. Transparency, calibrated trust, reliability [59] [10, 31, 60, 61, 62, 63, 64]
17 Trust in generative AI can be divided into cognitive trust and affective trust. Cognitive trust is based on a user’s evaluation of the AI platform’s ability to deliver accurate and reliable information, while affective trust stems from the user’s belief that the AI system considers their interests and well-being. Reliable, credible [65] [66, 67, 68]
18 The author actually criticizes the term trustworthy and ethical AI because they can humanize AI. Responsible AI [69] -
@@
In Table II, we provide a structured overview of various definitions and conceptualizations of trust and trustworthiness in AI, as discussed in the literature. This compilation allows for a comparative analysis of how trust in AI is framed across different domains, emphasizing dimensions such as ability, reliability, integrity, explainability and ethical alignment. From the initial 68 analysed papers, 18 papers are explicitly or implicitly define trust/trustworthiness in LLM or genAI. That is, 26.47% against 73.53% of papers that are not providing useful information for this task. 50 papers were not taken into consideration because they are either closed access, on the use of LLM to improve trust in a different domain e.g., radiology, measuring trust level of a certain group when using certain LLM system, or simply with a different focus.
@@
We found a tendency in reviewed studies in utilizing Mayer et al.’s work – “An Integrative Model of Organizational Trust”, published in 1995 – to build on definitions for what is trustworthiness in LLM – 33,33%, one third, of definitions (6/18 entries: 5 [30], 6 [40], 8 [43], 12 [52], 15 [58], 16 [59]) explicitly draw from Mayer et al.’s [10] organizational trust framework. Trust is described as a three factor construct consisting of ability, benevolence, and integrity [10]. Martell et al. [59] add predictability on Mayer et al.’s work, naming this as ABI+ factors of perceived trustworthiness towards AI:
  1. 1.
@@
Trust in AI, particularly in situations of uncertainty and vulnerability, is described as the belief that an AI agent will support an individual in achieving his or her goals [30]. Trust is fundamentally linked to vulnerability and uncertainty, as highlighted by multiple perspectives.
Hoy and Tschannen-Moran [50] define trust as a willingness to be vulnerable based on confidence in another party’s benevolence, reliability, competence, honesty, and openness. Mayer et al. [10] reinforce this by framing trust as a positive expectation of another’s intentions and behavior, emphasizing the risk involved in relying on an entity. Similarly, Lee and See [31] describe trust as the belief that an agent will assist in achieving one’s goals despite uncertainty and vulnerability, underscoring that trust in AI systems depends not just on their performance but also on users’ confidence in their intent and reliability. In this context, users assume the vulnerable position of relying on genAI models or proprietary platforms with the expectation that this risk will be beneficial to them, i.e. that it will help them achieve their goals despite possible misinformation or bias [52].
In the paper “Mitigative Strategies for Recovering From Large Language Model Trust Violations” [59], the authors present three different levels of trust calibration, as shown in Table III.
TABLE III: Types of Trust Calibration in Human-AI Interaction [59]
@@
Appropriate (Calibrated) Trust: Alignment between an individual’s perceived trust level and the AI system’s actual capabilities and performance. [60]
Over-Trust: Occurs when users: Perceive the AI as more capable than it truly is Treat the technology as a fully competent teammate Over-rely on AI outputs despite known limitations [61]
Under-Trust: Occurs when users: Underestimate the AI system’s capabilities Over-rely on their own judgment Dismiss potentially valuable AI contributions [62]
@@
In Table IV, we present the frequency of key terms associated with trustworthiness in AI, reflecting the most commonly emphasized aspects in the literature. This Table was produced by analysing the synonyms to trustworthy AI presented in Table II.
TABLE IV: Frequency of Trustworthiness-Related Terms Extracted from Reviewed Papers (2019-2025)